Home | Community | Message Board

MushroomMan Mycology
This site includes paid links. Please support our sponsors.


Welcome to the Shroomery Message Board! You are experiencing a small sample of what the site has to offer. Please login or register to post messages and view our exclusive members-only content. You'll gain access to additional forums, file attachments, board customizations, encrypted private messages, and much more!

Jump to first unread post Pages: | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |  [ show all ]
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21193736 -

berdinwall said:
I just think its bs. I get yalls point but id bet most of you put your money on it not being heads again on the 10th flip



:ducklol:

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21193738 -

:classictroll:

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: psi]
    #21194699 -

To be honest, betting on heads is the wrong answer. If you got nine tails in a row, most likely that is pure chance and the next flip has a 50:50 chance of being tails again. But maybe there is something wrong with the coin that makes tails more likely, in which case you would want to bet tails.

Since the first case it wouldn't matter which one you picked and the second it would be best to bet tails... You should always bet tails or more generally bet which ever side has come up MORE often.

In no case would you want to bet heads after nine tails.


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Edited by koods (01/29/15 02:11 AM)

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21194918 -

berdinwall said:
alright...so say you're flipping a quarter 10 times

and you flip heads 9 times in a row....are the odds that the last flip will be tails greater because of the previous flips? or still 50/50?

Someone once told me it was a matter of statistics. does statistics break down to true math? or do you just add a factor of statistics into the original 50/50 probability?

is there a proven mathematical answer?




No.

there are only two possible outcomes therefore  1/2=.5
or 50%


berdinwall said:
ok if you had to bet your last $500 on heads or tails for the last flip, which would you choose?





Tails if i watched all 9 previous flips.
The probability of going 10/10 on head is.

.5^10 

the odds of what you explained are already at .0019
its not gonna happen.


but to answer your question more directly, every throw of the coin has the odds of .5 regardless of what happens before it.

         
The formula to calculate odds is { Σ P(Ei) }to infinity and i=1



so the probability of two consecutive dice throws of the same number is (1/6)(1/6)

Edited by elax420 (01/29/15 04:33 AM)

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: elax420]
    #21195171 -

elax420 said:
berdinwall said:
alright...so say you're flipping a quarter 10 times

and you flip heads 9 times in a row....are the odds that the last flip will be tails greater because of the previous flips? or still 50/50?

Someone once told me it was a matter of statistics. does statistics break down to true math? or do you just add a factor of statistics into the original 50/50 probability?

is there a proven mathematical answer?




No.

there are only two possible outcomes therefore  1/2=.5
or 50%


berdinwall said:
ok if you had to bet your last $500 on heads or tails for the last flip, which would you choose?





Tails if i watched all 9 previous flips.
The probability of going 10/10 on head is.

.5^10 

the odds of what you explained are already at .0019
its not gonna happen.


but to answer your question more directly, every throw of the coin has the odds of .5 regardless of what happens before it.

         
The formula to calculate odds is { Σ P(Ei) }to infinity and i=1



so the probability of two consecutive dice throws of the same number is (1/6)(1/6)



The probability of two dice throws having the same number is 1:6.  The probability of you picking a number and having it come up twice is 1:36.

The first situation, the first roll can be anything so the odds of anything are 1:1. The second roll must match whatever number came up on the first roll, that is 1:6.


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Edited by koods (01/29/15 08:31 AM)

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21195194 -

This is a harder probability question. What is the chance to get three dice to add up to 15 in a single roll. (Rolling all three dice at once, the three face values add up to 15)?


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21195201 -

That's not statistics it's probability theory, yes the chance of it landing on heads is still 50/50 after 9 times in a row.  Even if it landed 99 times in a row on head it's still 50/50.

anddddddd im done lol just took a look at the rest of this thread, a little too dweeby for me :lol:


--------------------
L'enfer est plein de bonnes volontés et désirs

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: Cognitive_Shift]
    #21195224 -

If it landed 99 times on the same side, there is something wrong with the coin :lol:


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: koods]
    #21195264 -

Yeah most likely a rigged coin then anything :lol:


--------------------
L'enfer est plein de bonnes volontés et désirs

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21195282 -

just guessing here but is it 1:216?


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21195293 -

ok maybe its 1:180? i just realized there is more than one way to achieve 15.


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: Algo]
    #21195364 -

There are ten ways, so 10:216

3,6,6
4,5,6
5,4,6
4,6,5
5,6,4
6,5,4
6,4,5
6,3,6
6,6,3
5,5,5


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21195389 -

koods said:
This is a harder probability question. What is the chance to get three dice to add up to 15 in a single roll. (Rolling all three dice at once, the three face values add up to 15)?



10 in 216 I think


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: zappaisgod]
    #21195392 -

Son of a bitch I swear I didn't look.


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21196721 -

koods said:
elax420 said:
berdinwall said:
alright...so say you're flipping a quarter 10 times

and you flip heads 9 times in a row....are the odds that the last flip will be tails greater because of the previous flips? or still 50/50?

Someone once told me it was a matter of statistics. does statistics break down to true math? or do you just add a factor of statistics into the original 50/50 probability?

is there a proven mathematical answer?




No.

there are only two possible outcomes therefore  1/2=.5
or 50%


berdinwall said:
ok if you had to bet your last $500 on heads or tails for the last flip, which would you choose?





Tails if i watched all 9 previous flips.
The probability of going 10/10 on head is.

.5^10 

the odds of what you explained are already at .0019
its not gonna happen.


but to answer your question more directly, every throw of the coin has the odds of .5 regardless of what happens before it.

         
The formula to calculate odds is { Σ P(Ei) }to infinity and i=1



so the probability of two consecutive dice throws of the same number is (1/6)(1/6)



The probability of two dice throws having the same number is 1:6.  The probability of you picking a number and having it come up twice is 1:36.

The first situation, the first roll can be anything so the odds of anything are 1:1. The second roll must match whatever number came up on the first roll, that is 1:6.





No, that is not how you write it at all. If you throw two dice at the same time the odds of them coming up the same are 1/6 but if you know how to read and understand what consecutively mean the odds are P(e1)(e2). Very fucking basic dude.


By the way 1:36  is the same thing as (1/6)(1/6) which you write as .0277





koods said:
There are ten ways, so 10:216

3,6,6
4,5,6
5,4,6
4,6,5
5,6,4
6,5,4
6,4,5
6,3,6
6,6,3
5,5,5



once again .0462 is definitely preferred but you can also write 5/108. You literally have to simplify. Have you ever taken math?
:facepalm:

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21196765 -

Being a good citizen Muhommad Akbar thought he would make everyone on passenger planes safer by taking his own bomb onboard every flight. Afterall, he thought, what are the odds of there being two bombs on a plane.  :um:


--------------------
"If the foundations be destroyed, what can the righteous do."-King Solomon

And there was war in heaven: Michael and his angels fought against the dragon; and the dragon fought and his angels,

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: elax420]
    #21196899 -

elax420 said:
koods said:
There are ten ways, so 10:216

3,6,6
4,5,6
5,4,6
4,6,5
5,6,4
6,5,4
6,4,5
6,3,6
6,6,3
5,5,5



once again .0462 is definitely preferred but you can also write 5/108. You literally have to simplify. Have you ever taken math?
:facepalm:



He provided the explanation.  Do you always have to be so obnoxious?


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: zappaisgod]
    #21196944 -

I'm always amused when condescending people are wrong.


--------------------
NotSheekle said
“if I believed she was 16 I would become unattracted to her”

Extras: Filter Print Post Top
Re: Math paradox, whats the answer?t [Re: koods]
    #21196954 -

Yeah, but he isn't wrong.  Just obnoxious


--------------------

Extras: Filter Print Post Top
Re: Math paradox, whats the answer? [Re: berdinwall]
    #21196984 -

berdinwall said:
alright...so say you're flipping a quarter 10 times

and you flip heads 9 times in a row....are the odds that the last flip will be tails greater because of the previous flips? or still 50/50?





the odds are always 50/50, the odds of us getting his by an asteroid are 50/50, either
it happens or it doesnt, the odds of a train falling from the sky and killing you are
50/50, either it happens or it doesnt

Extras: Filter Print Post Top
Jump to top Pages: | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 |  [ show all ]


Similar ThreadsPosterViewsRepliesLast post
* Math question HELP!
( 1 2 all )
Shiithead 650 24 05/05/21 01:38 AM
by Shiithead
* 5-Year-Old Cuffed, Arrested in Florida
( 1 2 all )
C20H25N3O 3,948 36 03/30/05 09:25 PM
by C20H25N3O
* urgent math problem i need help getting solved
( 1 2 all )
phEight 2,917 35 04/01/05 04:27 PM
by freddurgan
* DVD audio in 5.1 ToTheSummit 813 2 03/14/03 04:04 PM
by Ulysees
* Rock climbers click here! Tony lamiche sends "the fly" 5.14d
( 1 2 3 all )
ummikko 5,067 52 03/08/04 05:10 PM
by ummikko
* CHALLENGE: Givin up 5 things for 23 days..Ooooh MAN
( 1 2 all )
ShroomFan 1,692 30 12/15/04 04:38 AM
by ShroomFan
* top 5 fav albums..
( 1 2 all )
BrAiN 3,647 27 01/19/05 07:29 PM
by drowneduck
* The story of how I turned $5 into $90 with no work flowstone 1,284 19 01/28/05 04:54 PM
by dressel11

Extra information
You cannot start new topics / You cannot reply to topics
HTML is disabled / BBCode is enabled
Moderator: Entire Staff
7,879 topic views. 10 members, 185 guests and 279 web crawlers are browsing this forum.
[ Show Images Only | Sort by Score | Print Topic ]
Search this thread:

Copyright 1997-2026 Mind Media. Some rights reserved.

Generated in 0.031 seconds spending 0.007 seconds on 16 queries.